# Geometry and factorization of multivariate Markov chains with applications to the swapping algorithm

### v4

@@ 2.4 $(C_{i})_{i=1}^{n}$-factorizable transition matrices with respect to a graph
In the literature of graphical model and Markov random field, factorization of probability masses with respect to the cliques of a graph has been investigated, and culminates in the Hammersley-Clifford theorem, see for instance (Wainwright, 2019, Chapter $11$). In this vein, in this subsection we consider the problem of factorizability of a transition matrix with respect to the cliques of a graph.
@@
Figure 2: Illustrations of $C_{1},C_{2},S_{1},S_{2}$ on a given $5$-node graph with $d=5$. We take $C_{1}=\{1,2,3\}$, $C_{2}=\{3,4,5\}$, $S_{1}=\{1,2,3\}$ and $S_{2}=\{4,5\}$. Note that these sets are all cliques of the graph and $S_{1}$,$S_{2}$ together form a partition of $\llbracket 5\rrbracket$.
@@ 2.5 Comparisons of mixing and hitting time parameters between $P$ and its information projections
Let $S\subseteq\llbracket d\rrbracket$. In Section 2.3, we have seen that $P^{(S)}\otimes P^{(-S)}$, the tensor product of the keep-$S$-in and leave-$S$-out transition matrix of a given $\pi$-stationary $P\in\mathcal{L}(\mathcal{X})$, arises naturally as an information projection of $P$ onto the space $\mathcal{L}_{S\otimes\llbracket d\rrbracket\backslash S}(\mathcal{X})$. The objective of this subsection is to investigate the relationship of hitting and mixing time parameters such as commute time, right spectral gap, log-Sobolev constant and Cheeger’s constant between $P$ and its information projections. These parameters play important roles in bounding the hitting or mixing time of $P$, see for instance Levin and Peres (2017); Saloff-Coste (1997); Montenegro and Tetali (2006); Aldous and Fill (2002) and the references therein.
@@
  • (Cheeger’s constant) Let $\emptyset\neq A\subseteq\mathcal{X}=\bigtimes_{i=1}^{d}\mathcal{X}^{(i)}$ and $A^{c}:=\mathcal{X}\backslash A$. We define
@@
    and hence the Cheeger’s constant of $P$ (Levin and Peres, 2017, Chapter $7.2$) is defined to be
@@
    Analogously, we define the Cheeger’s constant of the keep-$S$-in transition matrix $P^{(S)}$. Let $\emptyset\neq A^{(S)}\subseteq\mathcal{X}^{(S)}$ and $A^{(S)c}=\mathcal{X}^{(S)}\backslash A^{(S)}$. We then have
@@
    $\mathbb{E}_{x}(\tau_{y}(P))+\mathbb{E}_{y}(\tau_{x}(P))$ $=\sup_{0\leq f\leq 1;~{}f(x)=1,f(y)=0}\dfrac{1}{\mathcal{D}_{\pi}(f,f)}.$ (11)
@@
The main result of this subsection recalls a contraction principle in (Aldous and Fill, 2002, Proposition $4.44$): the hitting and mixing time parameters such as $\gamma,\alpha,\Phi,t_{c},t_{av}$ are at least “faster” for $P^{(S)}$ than the original chain $P$. We also establish new monotonicity results for the parameters $S\mapsto\mathbb{I}^{\pi^{(S)}}(P^{(S)})$ and $S\mapsto D_{KL}^{\pi^{(S)}}(P^{(S)}\|\Pi^{(S)})$ via the partition lemma presented in Theorem 2.4.
@@
where $\Pi\in\mathcal{L}(\pi)$ is the transition matrix with each row being $\pi$. All the above equalities hold if $T=\llbracket d\rrbracket$. The last two inequalities also hold for general $\pi$-stationary $P$.
@@
We note that it is perhaps possible to obtain tighter bounds in Corollary 2.3 by considering the corresponding hitting or mixing time parameters of the “restriction chains” as in Jerrum et al. (2004), see Section 3.1.
Proof of Corollary 2.3.
Once we have obtained the inequalities governing $P$ and $P^{(S)}$, we replace $P$ by $P^{(S)}$ and note that $\left(P^{(S)}\right)^{(T)}=P^{(T)}$ to reach at the inequalities governing $P^{(S)}$ and $P^{(T)}$. As a result, it suffices to prove only the inequalities between $P$ and $P^{(S)}$.
The case of $\gamma,\alpha,\Phi,t_{c},t_{av}$ have already been analyzed in (Aldous and Fill, 2002, Proposition $4.44$).
@@
The next result compares $\gamma(P)$ and $\gamma(P^{(S)}\otimes P^{(-S)})$ to obtain
@@
Let $\emptyset\neq S\subseteq\llbracket d\rrbracket$. Let $P\in\mathcal{L}(\pi)$ be a lazy, ergodic and $\pi$-reversible transition matrix. We have
@@
Proof.
As $P$ is lazy, both $P^{(S)},P^{(-S)}$ are also lazy by Proposition 2.3, and hence $\lambda_{2}(P^{(S)}\otimes P^{(-S)})=\max\{\lambda_{2}(P^{(S)}),\lambda_{2}(P^{(-S)})\}$. The desired result follows from Proposition 2.3. ∎
@@ 2.6 Some submodular functions arising in the information theory of multivariate Markov chains
This subsection is devoted to prove that the mappings $S\mapsto H(P^{(S)})$ and $S\mapsto D(P\|P^{(S)}\otimes P^{(-S)})$ are submodular in $S$. These two properties are analogous to the counterpart properties of the Shannon entropy and the mutual information of random variables: they are respectively monotonically non-decreasing submodular and submodular functions (see for example (Polyanskiy and Wu, 2022, Chapter $1.4,1.5$)). Note that in the i.i.d. case, the submodularity of Shannon entropy implies the Han’s inequality via the notion of self-bounding function, see for example Sason (2022, Corollary 2).
@@
Proof.
[…]
@@
Finally, we prove item (3). The monotonicity is shown in Proposition 2.3. We also note that
@@
which is a sum of a modular function and a supermodular function $-H(P^{(S)})$. As the sum of two supermodular functions is supermodular, $\mathbb{I}(P^{(S)})$ is supermodular.
@@ 3 Applications of projection chains to the design of MCMC samplers
We stress that the technique of projection chains is not limited to the swapping algorithm. The technique is broadly applicable to speedup mixing as long as we have a reversible multivariate Markov chains (such as particle-based MCMC) with stationary distribution $\pi$ and we are only interested in sampling from $\pi^{(S)}$, under the assumption that certain conditional distributions can be sampled from (recall Remark 2.5).
@@ 3.1 An improved projection sampler based on the swapping algorithm and its mixing time analysis
We first consider the special case of $d=2$-temperature swapping algorithm. The main results of this section (Corollary 3.2 and 3.3 below) state that various mixing time parameters of the keep-$\{2\}$-in (or leave-$\{1\}$-out) chain based on a two-temperature swapping algorithm are improved by at least a factor of the dimension of the underlying state space than the original swapping algorithm. We also offer an intuitive explanation on why this projection sampler accelerates over the original algorithm in Remark 3.1.
To this end, let us fix the notations and briefly recall the dynamics of the swapping algorithm. Much of our exposition in this example follows that in Bhatnagar and Randall (2016). Let $P_{0}$ be an ergodic and reversible chain with stationary distribution being the discrete uniform $\pi_{0}$ on $\mathcal{X}$. Denote the Boltzmann-Gibbs distribution associated with energy function $\mathcal{H}:\mathcal{X}\to\mathbb{R}$ at inverse temperature $\beta\geq 0$ to be
@@
At each step, the swapping chain chooses an inverse temperature $\beta_{i}$ uniformly at random then chooses uniformly at random between the following two moves:
@@
In the special case of $d=2$, the swapping algorithm amounts to running two Markov chains (aka particles) simultaneously with one at inverse temperature 00 and the other one at the target $\beta$ coupled with swapping moves between the states of these two chains. With these choices, $\pi_{sw}=\pi_{0}\otimes\pi_{\beta}$, the product distribution of $\pi_{0}$ and $\pi_{\beta}$. The keep-$\{2\}$-in (or $2$nd marginal as in Definition 2.3) Markov chain with transition matrix $P_{sw}^{(2)}$ can be written as, for $x^{2},y^{2}\in\mathcal{X}$,
@@
In the terminologies of Jerrum et al. (2004), the projection chain (i.e. the notation $\overline{P}$ therein) of $P_{sw}$ with respect to the above partition is $P_{sw}^{(2)}$, while the restriction chains (i.e. the notation $P_{i}$ therein) on the state space $\Omega_{x}$ can be shown to be, for each $x\in\mathcal{X}$,
@@
$\gamma((P_{sw})_{x})=\dfrac{1}{4}\gamma(P_{0}),\quad\alpha((P_{sw})_{x})=\dfrac{1}{4}\alpha(P_{0}).$ (12)
Since the projection chain $P_{sw}^{(2)}$ and the restriction chains $(P_{sw})_{x}$ are ergodic, Theorem $1$ and Theorem $4$ in Jerrum et al. (2004) can be applied in this setting to yield
@@
In view of the above result, it is thus advantageous to use the projection sampler $P_{sw}^{(2)}$ rather than the original $P_{sw}$ to sample from $\pi_{\beta}$, as the former is at least three times faster than the latter in terms of relaxation or log-Sobolev time. Note that this result is also an improvement towards the general contraction principle presented in Proposition 2.3.
Let us now specialize into $\mathcal{X}=\{0,1\}^{N}$ with $N\in\mathbb{N}$, and consider $P_{0}$ being the transition matrix of the simple random walk on the hypercube $\mathcal{X}$. In other words, at each time a coordinate is chosen uniformly at random, and the entry of the chosen coordinate is flipped to the other with probability $1$ while keeping all other coordinates unchanged. It is well-known (see e.g. (Jerrum et al., 2004, Section $4.5$)) that $\alpha(P_{0})=\gamma(P_{0})/2=\frac{1}{N+1}$, and hence, in view of (12), we arrive at
$\gamma((P_{sw})_{x})=\dfrac{1}{2}\dfrac{1}{N+1},\quad\alpha((P_{sw})_{x})=\dfrac{1}{4}\dfrac{1}{N+1}.$ (13)
Denote the parameter $\Gamma$ (i.e. the notation $\gamma$ in Jerrum et al. (2004)) to be
$\Gamma:=\max_{x\in\mathcal{X}}\max_{y\in\mathcal{X}}\left(1-\sum_{z\in\mathcal{X}}P_{sw}((y,x),(z,x))\right).$ (14)
@@
Using again Theorem $1$ and Theorem $4$ in Jerrum et al. (2004) leads to
@@
A celebrated result (Diaconis and Saloff-Coste, 1996, Page $697$) gives that
$\dfrac{1}{2\alpha(P)}\leq T_{mix}(P,1/e)\leq\dfrac{4+\log\log(1/\min_{x}\pi(x))}{\alpha(P)}.$ (15)
@@
Many models in statistical physics satisfy a polynomial in $N$ oscillation of $\mathcal{H}$ with $\mathrm{Osc}(\mathcal{H})=\mathcal{O}(N^{k})$ for some positive integers $k$, for instance the Curie-Weiss model on a complete graph or the Ising model on finite grid (Nardi and Zocca, 2019). From the viewpoint of MCMC again, the above results indicate that at times it is advantageous to simulate the keep-$\{2\}$-in $P_{sw}^{(2)}$ over $P_{sw}$ with a speedup of at least a factor of $N/\log(\beta N)$.

### v5

@@ 2.4 $(C_{i})_{i=1}^{n}$-factorizable transition matrices with respect to a graph
In the literature of graphical model and Markov random field, factorization of probability masses with respect to the cliques of a graph has been investigated, and culminates in the Hammersley-Clifford theorem, see for instance [42, Chapter $11$]. In this vein, in this subsection we consider the problem of factorizability of a transition matrix with respect to the cliques of a graph.
@@
Figure 3: Illustrations of $C_{1},C_{2},S_{1},S_{2}$ on a given 55-node graph with $d=5$. We take $C_{1}=\{1,2,3\}$, $C_{2}=\{3,4,5\}$, $S_{1}=\{1,2,3\}$ and $S_{2}=\{4,5\}$. Note that these sets are all cliques of the graph and $S_{1}$$S_{2}$ together form a partition of $\llbracket 5\rrbracket$.
@@ 2.5 Comparisons of mixing and hitting time parameters between $P$ and its information projections
Let $S\subseteq\llbracket d\rrbracket$. In Section 2.3, we have seen that $P^{(S)}\otimes P^{(-S)}$, the tensor product of the keep-$S$-in and leave-$S$-out transition matrix of a given $\pi$-stationary $P\in\mathcal{L}(\mathcal{X})$, arises naturally as an information projection of $P$ onto the space $\mathcal{L}_{S\otimes\llbracket d\rrbracket\backslash S}(\mathcal{X})$. The objective of this subsection is to investigate the relationship of hitting and mixing time parameters such as commute time, right spectral gap, log-Sobolev constant and Cheeger’s constant between $P$ and its information projections. These parameters play important roles in bounding the hitting or mixing time of $P$, see for instance [25, 38, 32, 1] and the references therein.
@@
  • (Cheeger’s constant) Let $\emptyset\neq A\subseteq\mathcal{X}=\bigtimes_{i=1}^{d}\mathcal{X}^{(i)}$ and $\AA^{c}:=\mathcal{X}\backslash A$. We define
@@
    and hence the Cheeger’s constant of $P$ [25, Chapter $7.2$] is defined to be
@@
[…]
@@
    (12) $\mathbb{E}_{x}(\tau_{y}(P))+\mathbb{E}_{y}(\tau_{x}(P))$ $=\sup_{0\leq f\leq 1;~f(x)=1,f(y)=0}\dfrac{1}{\mathcal{D}_{\pi}(f,f)}.$
@@
The main result of this subsection recalls a contraction principle in [1, Proposition $4.44$]: the hitting and mixing time parameters such as $\gamma,\alpha,\Phi,t_{c},t_{av}$ are at least “faster” for $P^{(S)}$ than the original chain $P$. We also establish new monotonicity results for the parameters $S\mapsto\mathbb{I}^{\pi^{(S)}}(P^{(S)})$ and $S\mapsto D_{KL}^{\pi^{(S)}}(P^{(S)}\|\Pi^{(S)})$ via the partition lemma presented in Theorem 2.21.
@@
where $\Pi\in\mathcal{L}(\pi)$ is the transition matrix with each row being $\pi$. All the above equalities hold if $T=\llbracket d\rrbracket$. The last two inequalities also hold for general $\pi$-stationary $P$ without reversibility.
@@
We note that it is perhaps possible to obtain tighter bounds in Corollary 2.39 by considering the corresponding hitting or mixing time parameters of the “restriction chains” as in [23], see Section 3.1.
Proof 2.41 (Proof of Corollary 2.39).
Once we have obtained the inequalities governing $P$ and $P^{(S)}$, we replace $P$ by $P^{(S)}$ and note that $\left(P^{(S)}\right)^{(T)}=P^{(T)}$ to reach at the inequalities governing $P^{(S)}$ and $P^{(T)}$. By recalling Remark 2.19, $P^{(S)}$ can be written as
$P^{(S)}(x^{(S)},y^{(S)})=\sum_{y^{(-S)}}\mathbb{E}_{x^{(-S)}\sim\pi(\cdot|x^{(S)})}\left[P\left((x^{(S)},x^{(-S)}),(y^{(S)},y^{(-S)})\right)\right],$
hence
$\quad\left(P^{(S)}\right)^{(T)}(x^{(T)},y^{(T)})$
$=\sum_{y^{(S\setminus T)}}\mathbb{E}_{x^{(S\setminus T)}\sim\pi^{(S)}(\cdot|x^{(T)})}\left[P^{(S)}\left((x^{(T)},x^{(S\setminus T)}),(y^{(T)},y^{(S\setminus T)})\right)\right]$
$=\sum_{y^{(S\setminus T)}}\mathbb{E}_{x^{(S\setminus T)}\sim\pi^{(S)}(\cdot|x^{(T)})}\left[\sum_{y^{(-S)}}\mathbb{E}_{x^{(-S)}\sim\pi(\cdot|x^{(S)})}\left[P\left((x^{(T)},x^{(S\setminus T)},x^{(-S)}),(y^{(T)},y^{(S\setminus T)},y^{(-S)})\right)\right]\right]$
$=\sum_{y^{(-T)}}\mathbb{E}_{x^{(-T)}\sim\pi(\cdot|x^{(T)})}\left[P\left((x^{(T)},x^{(-T)}),(y^{(T)},y^{(-T)})\right)\right]$
$=P^{(T)}(x^{(T)},y^{(T)}).$
As a result, it suffices to prove only the inequalities between $P$ and $P^{(S)}$.
The case of $\gamma,\alpha,\Phi,t_{c},t_{av}$ have already been analyzed in [1, Proposition $4.44$].
@@
For non-reversible Markov chains, these above quantities can be far from sharply bounding mixing times. A good alternative is multiplicative spectral gap [32, Chapter 1], which is the gap between $1$ and second largest singular value. For a finite Markov chain with transition matrix $P$, the multiplicative spectral gap is defined as
$\gamma_{M}(P):=\gamma(\sqrt{PP^{*}})=1-\left\|P\right\|_{\ell_{0}^{2}(\pi)\to\ell_{0}^{2}(\pi)},$
where
$\left\|P\right\|_{\ell_{0}^{2}(\pi)\to\ell_{0}^{2}(\pi)}:=\sup_{f\in\ell_{0}^{2}(\pi)}\frac{\left\|Pf\right\|_{2}}{\left\|f\right\|_{2}},\quad\ell_{0}^{2}(\pi)=\left\{f\in\ell^{2}(\pi):\pi(f)=0\right\}.$
Next, we provide the contraction principle based on the multiplicative spectral gap for non-reversible chains.
@@
Let $\emptyset\neq T\subseteq S\subseteq\llbracket d\rrbracket$ and $P$ be an ergodic transition matrix. Without assuming the reversibility of $P$, we have
$\gamma_{M}(P)\leq\gamma_{M}(P^{(S)})\leq\gamma_{M}(P^{(T)}).$
Proof 2.43.
It suffices to prove the first inequality. According to Remark 2.19, the projected chain $P^{(S)}$ can be used to characterize the following movement on $\mathcal{X}^{(S)}$:
  1. (i)
    Starting from $x^{(S)}$, draw $x^{(-S)}\sim\pi(\cdot|x^{(S)})$;
  2. (ii)
    Draw $(y^{(S)},y^{(-S)})\sim P\left((x^{(S)},x^{(-S)}),\cdot\right)$;
  3. (iii)
    Update $x^{(S)}\leftarrow y^{(S)}$.
Next, we define $K_{S}:\ell^{2}(\pi)\to\ell^{2}(\pi^{(S)})$ and $J_{S}:\ell^{2}(\pi^{(S)})\to\ell^{2}(\pi)$ as
$K_{S}f(x^{(S)}):=\mathbb{E}_{x^{(-S)}\sim\pi(\cdot|x^{(S)})}\left[f(x^{(S)},x^{(-S)})\right],\quad f\in\ell^{2}(\pi),$
$J_{S}g(x^{(S)},x^{(-S)}):=g(x^{(S)}),\quad g\in\ell^{2}(\pi^{(S)}),$
then it is easy to see that $P^{(S)}=K_{S}PJ_{S}$. Moreover, $K_{S}$ and $J_{S}$ are adjoint operators, since for any $f\in\ell^{2}(\pi)$ and $g\in\ell^{2}(\pi^{(S)})$,
$\left\langle K_{S}f,g\right\rangle_{\pi^{(S)}}$ $=\sum_{x^{(S)}}g(x^{(S)})\pi^{(S)}(x^{(S)})\sum_{x^{(-S)}}f(x^{(S)},x^{(-S)})\pi(x^{(-S)}|x^{(S)})$
$=\sum_{x}f(x)g(x^{(S)})\pi(x)$
$=\left\langle f,J_{S}g\right\rangle_{\pi}.$
Observing that $J_{S}$ is an isometric embedding with $\left\|J_{S}\right\|_{\ell_{0}^{2}(\pi^{(S)})\to\ell_{0}^{2}(\pi)}=1$, we have
$\left\|P^{(S)}\right\|_{\ell_{0}^{2}(\pi^{(S)})\to\ell_{0}^{2}(\pi^{(S)})}$ $=\left\|K_{S}PJ_{S}\right\|_{\ell_{0}^{2}(\pi^{(S)})\to\ell_{0}^{2}(\pi^{(S)})}$
$\leq\left\|K_{S}\right\|_{\ell_{0}^{2}(\pi)\to\ell_{0}^{2}(\pi^{(S)})}\cdot\left\|P\right\|_{\ell_{0}^{2}(\pi)\to\ell_{0}^{2}(\pi)}\cdot\left\|J_{S}\right\|_{\ell_{0}^{2}(\pi^{(S)})\to\ell_{0}^{2}(\pi)}$
$=\left\|J_{S}\right\|_{\ell_{0}^{2}(\pi^{(S)})\to\ell_{0}^{2}(\pi)}\cdot\left\|P\right\|_{\ell_{0}^{2}(\pi)\to\ell_{0}^{2}(\pi)}\cdot\left\|J_{S}\right\|_{\ell_{0}^{2}(\pi^{(S)})\to\ell_{0}^{2}(\pi)}$
$=\left\|P\right\|_{\ell_{0}^{2}(\pi)\to\ell_{0}^{2}(\pi)},$
where we have used the fact that adjoint operators share the same norm. Then the result follows.
The next result compares $\gamma_{M}(P)$ and $\gamma_{M}(P^{(S)}\otimes P^{(-S)})$ to obtain
Corollary 2.44.
Let $\emptyset\neq S\subseteq\llbracket d\rrbracket$. Let $P$ be an ergodic and $\pi$-stationary transition matrix. We have
$\gamma_{M}(P)\leq\gamma_{M}(P^{(S)}\otimes P^{(-S)}).$
If we further assume that $P$ is lazy and reversible, then we have
@@
Proof 2.45.
It is easy to verify that $\left(P^{(S)}\otimes P^{(-S)}\right)^{*}=\left(P^{(S)}\right)^{*}\otimes\left(P^{(-S)}\right)^{*}$ and
$\left(P^{(S)}\otimes P^{(-S)}\right)\left(P^{(S)}\otimes P^{(-S)}\right)^{*}=P^{(S)}\left(P^{(S)}\right)^{*}\otimes P^{(-S)}\left(P^{(-S)}\right)^{*}.$
Recalling for two reversible transition matrices $Q_{1}\in\mathcal{L}(\pi_{1})$ and $Q_{2}\in\mathcal{L}(\pi_{2})$ with non-negative eigenvalues, it is well known that
$\lambda_{2}(Q_{1}\otimes Q_{2})=\max\left\{\lambda_{2}(Q_{1}),\lambda_{2}(Q_{2})\right\},$
then the first result comes from Corollary 2.42. If $P$ is further assumed to be lazy and reversible, both $P^{(S)},P^{(-S)}$ are also lazy by Proposition 2.20, and hence $\lambda_{2}(P^{(S)}\otimes P^{(-S)})=\max\{\lambda_{2}(P^{(S)}),\lambda_{2}(P^{(-S)})\}$. The desired result follows from Corollary 2.39.
@@ 2.6 Some submodular functions arising in the information theory of multivariate Markov chains
This subsection is devoted to prove that the mappings $S\mapsto H(P^{(S)})$ and $S\mapsto D(P\|P^{(S)}\otimes P^{(-S)})$ are submodular in $S$. These two properties are analogous to the counterpart properties of the Shannon entropy and the mutual information of random variables: they are respectively monotonically non-decreasing submodular and submodular functions (see for example [36, Chapter $1.4,1.5$]). Note that in the i.i.d. case, the submodularity of Shannon entropy implies the Han’s inequality via the notion of self-bounding function, see for example [39, Corollary 2].
@@
Proof 2.47.
We first prove item (1). Let $S\subseteq T\subseteq\llbracket d\rrbracket$ and suppose $i\in\llbracket d\rrbracket\backslash T$. By the definition of submodularity, it suffices to show that
@@
Finally, we prove item (3). The monotonicity is shown in Corollary 2.39. We also note that
@@
which is a sum of a modular function and a supermodular function $H(P^{(S)})$. As the sum of two supermodular functions is supermodular, $\mathbb{I}(P^{(S)})$ is supermodular.
@@ 3 Applications of projection chains to the design of MCMC samplers
We stress that the technique of projection chains is not limited to the swapping algorithm. In view of Corollary 2.39, the technique is broadly applicable to speedup mixing of multivariate Markov chains (such as particle-based MCMC) with stationary distribution $\pi$ and we are only interested in sampling from $\pi^{(S)}$, under the assumption that certain conditional distributions can be sampled from (recall Remark 2.19). In this section, we focus on a particular reversible multivariate Markov chain (swapping algorithm), as finer or more precise mixing time results can usually be obtained for reversible chains.
@@ 3.1 An improved projection sampler based on the swapping algorithm and its mixing time analysis
We first consider the special case of $d=2$-temperature swapping algorithm. The main results of this section (Corollaries 3.2 and 3.3 below) state that various mixing time parameters of the keep-$\{2\}$-in (or leave-$\{1\}$-out) chain based on a two-temperature swapping algorithm are improved by at least a factor of the dimension of the underlying state space than the original swapping algorithm. We also offer an intuitive explanation on why this projection sampler accelerates over the original algorithm in Remark 3.5.
To this end, let us fix the notations and briefly recall the dynamics of the swapping algorithm. Much of our exposition in this example follows that in [2]. Let $P_{0}$ be an ergodic and reversible chain with stationary distribution being the discrete uniform $\pi_{0}$ on $\mathcal{X}$. Denote the Boltzmann-Gibbs distribution associated with energy function $\mathcal{H}:\mathcal{X}\to\mathbb{R}$ at inverse temperature $\beta\geq 0$ to be
@@
At each step, the swapping chain chooses uniformly at random between the following two moves:
@@
In the special case of $d=2$, the swapping algorithm amounts to running two Markov chains (aka particles) simultaneously with one at inverse temperature 0 and the other one at the target $\beta$ coupled with swapping moves between the states of these two chains. With these choices, $\pi_{sw}=\pi_{0}\otimes\pi_{\beta}$, the product distribution of $\pi_{0}$ and $\pi_{\beta}$. The keep-$\{2\}$-in (or 22nd marginal as in Definition 2.10) Markov chain with transition matrix $P_{sw}^{(2)}$ can be written as, for $x^{2},y^{2}\in\mathcal{X}$,
@@
In the terminologies of [23], the projection chain (i.e. the notation $\overline{P}$ therein) of $P_{sw}$ with respect to the above partition is $P_{sw}^{(2)}$, while the restriction chains (i.e. the notation $P_{i}$ therein) on the state space $\Omega_{x}$ can be shown to be, for each $x\in\mathcal{X}$,
@@
(13) $\gamma((P_{sw})_{x})=\dfrac{1}{4}\gamma(P_{0}),\quad\alpha((P_{sw})_{x})=\dfrac{1}{4}\alpha(P_{0}).$
Since the projection chain $P_{sw}^{(2)}$ and the restriction chains $(P_{sw})_{x}$ are ergodic, Theorem $1$ and Theorem $4$ in [23] can be applied in this setting to yield
@@
In view of the above result, it is thus advantageous to use the projection sampler $P_{sw}^{(2)}$ rather than the original $P_{sw}$ to sample from $\pi_{\beta}$, as the former is at least three times faster than the latter in terms of relaxation or log-Sobolev time. Note that this result is also an improvement towards the general contraction principle presented in Corollary 2.39.
Let us now specialize into $\mathcal{X}=\{0,1\}^{N}$ with $N\in\mathbb{N}$, and consider $P_{0}$ being the transition matrix of the simple random walk on the hypercube $\mathcal{X}$. In other words, at each time a coordinate is chosen uniformly at random, and the entry of the chosen coordinate is flipped to the other with probability $1$ while keeping all other coordinates unchanged. It is well-known (see e.g. [23, Section $4.5$]) that $\alpha(P_{0})=\gamma(P_{0})/2=\frac{1}{N+1}$, and hence, in view of (13), we arrive at
(14) $\gamma((P_{sw})_{x})=\dfrac{1}{2}\dfrac{1}{N+1},\quad\alpha((P_{sw})_{x})=\dfrac{1}{4}\dfrac{1}{N+1}.$
Denote the parameter $\Gamma$ (i.e. the notation $\gamma$ in [23]) to be
(15) $\Gamma:=\max_{x\in\mathcal{X}}\max_{y\in\mathcal{X}}\left(1-\sum_{z\in\mathcal{X}}P_{sw}((y,x),(z,x))\right).$
@@
Using again Theorem $1$ and Theorem $4$ in [23] leads to
@@
A celebrated result [13, Page $697$] gives that
(16) $\dfrac{1}{2\alpha(P)}\leq T_{mix}(P,1/e)\leq\dfrac{4+\log\log(1/\min_{x}\pi(x))}{\alpha(P)}.$
@@
Many models in statistical physics satisfy a polynomial in $N$ oscillation of $\mathcal{H}$ with $\mathrm{Osc}(\mathcal{H})=\mathcal{O}(N^{k})$ for some positive integers $k$, for instance the Curie-Weiss model on a complete graph or the Ising model on finite grid [33]. From the viewpoint of MCMC again, the above results indicate that at times it is advantageous to simulate the keep-$\{2\}$-in $P_{sw}^{(2)}$ over $P_{sw}$ with a speedup of at least a factor of $N/\log(\beta N)$.